OpenStudy (anonymous):

In how many ways can the letters in the word PROPORTIONALITY be arranged so that the vowels and consonants still occupy the same place?

OpenStudy (anonymous):

Let's try : Consonants: 8 places, 6 different letters => 6! Vowels: 7 places, 4 different letters => 4! 4!*6!=17280 ? But I'm quite sure I'm wrong :D

OpenStudy (anonymous):

well the answer is 2, 721, 600

OpenStudy (anonymous):

but i don't know how to get it

OpenStudy (anonymous):

^^ haha...

OpenStudy (anonymous):

are we considering Y a vowel? lol

OpenStudy (anonymous):

lol:) u were close

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

Really ? It's not a vowel in english ??

OpenStudy (anonymous):

its just a,e,i,o,u isn't it

OpenStudy (anonymous):

ok, im counting 9 consonants, so that makes 9!, but there are 2 P's, 2 R's, and 2 T's, so the number of arrangements is: \[\frac{9!}{2!2!2!} = 45360\] Now looking at the vowels, there are 6 of them, which is 6!, but there are 3 O's, and 2 I's, so the number of arrangements is: \[\frac{6!}{3!2!} = 60\] 45360*60 = 2721600 total combinations.

OpenStudy (anonymous):

Thanks for the explanation (and the discovery of Y not beeing a vowel :D)

OpenStudy (anonymous):

oh k thanks for the working:)

OpenStudy (phi):

Hmm, this is a bit bogus for a question. The y is clearly acting as a vowel.

OpenStudy (anonymous):

i never thought y was a vowel until i looked it up on wikipedia lol

OpenStudy (anonymous):

ive always seen it as not a vowel, but sometimes you hear, "AEIOU and sometimes Y" lol

OpenStudy (anonymous):

yeah since primary school we've only been taught a,e,i,o,u

OpenStudy (anonymous):

That's why I keep thinking that english pronunciation is hard >.< Always changing...

OpenStudy (anonymous):

which country are you from?

OpenStudy (anonymous):

France (I know, we suck at foreign language ^^)

OpenStudy (anonymous):

i don't know about that lol. france has good music:)

OpenStudy (anonymous):

Maybe ^^ Wondering what kind of french music can come to your ears

OpenStudy (anonymous):

Alors on danse:)

OpenStudy (anonymous):

Ok. Didn't know french music could travel that much apart from Daft Punk. Btw, I think Stromae ("alors on danse") is actually from Belgium

OpenStudy (anonymous):

he is from belgium, do they speak french in belgium?

OpenStudy (anonymous):

is daft punk french i didn't know that

OpenStudy (anonymous):

Daft punk is french yep. For Belgians it's a story by itself... 50% speak french 50% dutch. There's kind of a political crisis currently there.

OpenStudy (anonymous):

i see lol. thats interesting to know:)